OpenStudy (anonymous):

what is 55/100 in simplest form

OpenStudy (anonymous):

11/20

OpenStudy (unklerhaukus):

\[\frac{55}{100}=\frac{5\times11}{5\times 20}=\frac{\cancel{5\times}11}{\cancel{5\times} 20}\]